Please help me solve this word problem...

Martina is currently 12 years older than her cousin Joey. In 4 years she will be 3 times as old as Joey. Use this information to answer the following questions.
If you let x represent Joey's current age, what expression can you use to represent Martina's current age?
If Joy's be x then Martin's be x+12 (because 12 years older than joy's age)

Found 2 solutions by macston, jhaashok630:
Answer by macston(5194)

x=Joey's age; m=martina's age=x+12
The only question asked is to express Martina's current age in terms of Joey's so
ANSWER:m=x+12 years
TRY TO FIND MARTINA'S AND JOEY'S AGES
The following included to just check your work if you decide to go on.
m+4=3(x+4)
m+4=3x+12 Substitute for m
x+12+4=3x+12
x+16=3x+12 Subtract x+12 from each side.
4=2x Divide each side by 2.
2=x Joey is 2 years old
m=x+12 years=2 years+12 years=14 years Martina is 14 years old.
CHECK
m+4 years=3(x+4 years)
14 years+4 years=3(2 years+4years)
18 years=3(6 years)
18 years=18 years

Answer by jhaashok630(18)

If Joy's be x then Martin's be x+12 (because 12 years older than joy's age)

In 4 years Joy's = x+4
In 4 years Martin's = x+12+4
In 4 years Martin's will be 3 times as old as Joey
x+12+4=3(x+4) Solve for x
x+16=3x+12
x=2
Current age of Joey 2